OpenStudy (deadshot):

What is the simplified form of 9 times y to the fifth power over 15 times x to the seventh power • 5 x cubed over 45 y squared ?

OpenStudy (deadshot):

\[\frac{ 9*y^5 }{ 15*x^7 }\times \frac{ 5x^3 }{ 45y^2 }\]

OpenStudy (anonymous):

\[{\cancel{9\times5}\over15\times\cancel{45}}\times{y^{5-2}\over x^{7-3}}=\frac{y^3}{15x^4} \]

OpenStudy (deadshot):

Thanks!